Using the separation of variables technique, derive the time-independent Schrodinger equation (T.I.S.E.) in three dimensions. 


1
Expert's answer
2020-07-21T13:43:01-0400

3-D Time-Independent Schrodinger Equation


(−ℏ2/2m)∇2ψ(r)+V(r)ψ(r)=Eψ(r)


Since we are dealing with a 3-dimensional figure, we need to add the 3 different axes into the Schrondinger equation:


−ℏ2/2m(d2ψ(r)/dx2+d2ψ(r)/dy2+d2ψ(r)/dz2)=Eψ(r)


The easiest way in solving this partial differential equation is by having the wavefunction equal to a product of individual function for each independent variable (e.g., the Separation of Variables technique):


ψ(x,y,z)=X(x)Y(y)Z(z)


Now each function has its own variable:


X(x) is a function for variable x only


Y(y) function of variable y only


Z(z) function of variable z only


Now substitute Equation 4 into Equation 3 and divide it by the product: xyz:


d2ψ/dx2=(YZ)d2X/dx2⇒(1/X)d2X/dx2


d2ψ/dy2=(XZ)d2Y/dy2⇒(1/Y)d2Y/dy2


d2ψ/dz2=(XY)d2Z/dz2⇒(1/Z)d2Z/dz2


E is an energy constant, and is the sum of x, y, and z. For this to work, each term must equal its own constant. For example,


(d2X/dx2(+(2m/ℏ2xX=0


Now separate each term in Equation to equal zero:


(d2X/dx2)+(2m/ℏ2xX=0


(d2Y/dy2)+(2m/ℏ2yY=0


(d2Z/dz2)+(2m/ℏ2zZ=0


Now we can add all the energies together to get the total energy:

Ex+Ey+Ez =E


Now the equations are very similar to a 1-D box and the boundary conditions are identical, i.e.,


n=1,2,..∞


Use the normalization wavefunction equation for each variable:


ψ(x)={(√2/Lx)sinnπxLx } if 0≤x≤L0if L<x<0


Normalization wavefunction equation for each variable


X(x)=(√2/Lx)sin(nxπx/Lx)


Y(y)=(√2/Ly)sin(nyπy/Ly)


Z(z)=(√2/Lz)sin(nzπz/Lz)


The limits of the three quantum numbers

nx=1,2,3,...∞

ny=1,2,3,...∞

nz=1,2,3,...∞

For each constant use the de Broglie Energy equation:

εx=nx2h2/8mLx2

with nx=1,2,3...∞

Do the same for variables ny and nz. Combine wave functions for all axis , So overall wavefunctions inside a 3D box is

ψ(r)=(√8/L)sin(nxπx/Lx)sin(nyπy/Ly)sin(nzπz/Lz)